OpenStudy (anonymous):

Help!! Indicate in standard form the equation of the line passing through the given points. P(6, 2), Q(8, -4)

OpenStudy (anonymous):

remember the genral equation of a line passing through two points \((x_1,y_1)\;{\rm and}\;(x_2,y_2)\) is given by\[ \frac{y-y_1}{y_1-y_2}=\frac{x-x_1}{x_1-x_2} \]

OpenStudy (anonymous):

plug in the co-ordinates of the two points in the equation

OpenStudy (anonymous):

Standard form requires the form: Ax+By=C The first thing we want to do is find the slope of this equation: Slope is rise/run, or change in y/change in x. From P to Q, x changes from 6 to 8, a total change of 2. Likewise, y changes by -6. Therefore, m=-6/2, or -3 simplified. The slope m will be the "A" from the formula stated above. The only thing we need to remember is that x is given positive priority. What this means is, you do what you must to make the coefficient for x positive (so A will actually be 3). So far we have: y=-3x+C To find C, we have to find the y intercept to turn this two variable equation into a single variable equation. At the y intercept, x=0. We can continue the pattern that the slope gives us to trace our way to that point: at x=6, y=2 as stated by the points given. Following this pattern, x=4 must give us y=8 (since for every -2 unit change in x, we have +6 unit change for y). Eventually we get x=0, y=20, our y intercept. The y intercept is also the C of our formula, giving us: y=-3x+20 This, when transformed into standard form, looks like: -3x-y+20 -3x-y=-20 3x+y=20 This can be checked by plugging in our points, which turn out to work. Hope this helps! If there is any more explanation I can give, just let me know.

OpenStudy (anonymous):

Cough